# Learning WENO for entropy stable schemes to solve conservation laws

### v1

@@ 3 Sign-Preserving WENO Reconstructions
A typical WENO reconstruction [23] produces a $(2k-1)$th-order accurate reconstruction by taking a convex combination of the $12k-$ 1 candidate polynomials considered in the $k$th-order ENO reconstruction. A third-order WENO has the following form for the left and right reconstructions of the variable $z$ at the interface $x_{i+\frac{1}{2}}$:
$z_{i+\frac{1}{2}}^{-}$ $=\hskip 9.95845ptw_{0,i+\frac{1}{2}}\left(\frac{z_{i}}{2}+\frac{z_{i+1}}{2}\right)+w_{1,i+\frac{1}{2}}\left(-\frac{z_{i-1}}{2}+\frac{3z_{i}}{2}\right),$ (14)
$z_{i+\frac{1}{2}}^{+}$ $=\tilde{w}_{0,i+\frac{1}{2}}\left(-\frac{z_{i+2}}{2}+\frac{3z_{i+1}}{2}\right)+\tilde{w}_{1,i+\frac{1}{2}}\left(\frac{z_{i}}{2}+\frac{z_{i+1}}{2}\right),$
where the weights must be chosen to: (i) ensure third-order accuracy of the reconstructions $z_{i+\frac{1}{2}}^{\pm}$ for smooth solutions and (ii) provide minimal weight to linear polynomials on stencils containing discontinuities. In general, WENO reconstructions do not satisfy the sign property, and their use in the TeCNO framework described by (13) does not yield entropy stable schemes. Thus, a WENO reconstruction must be specifically designed to also satisfy the sign property.
3.1 SP-WENO [16]
@@
$C_{1}(\theta_{i}^{+},\theta_{i+1}^{-})=\begin{cases}\frac{1}{8}\left(\frac{f^{+}}{(f^{+})^{2}+(f^{-})^{2}}\right)&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}<0,\psi^{+}\neq-1\\ 0&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}=-1\\ -\frac{3}{8}&\text{ if }\theta_{i}^{+}=1\text{ or }\psi^{+}\geq 0,|\theta_{i}^{+}|\leq 1\\ \frac{1}{8}&\text{ if }\psi^{+}\geq 0,|\theta_{i}^{+}|>1\end{cases},$
@@
$f^{+}(\theta_{i}^{+},\theta_{i+1}^{-}):=\begin{cases}\frac{1}{1+\psi^{+}}&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}\neq-1\\ 1&\text{ otherwise}\end{cases},\hskip 14.22636ptf^{-}(\theta_{i}^{+},\theta_{i+1}^{-}):=f^{+}(\theta_{i+1}^{-},\theta_{i}^{+}).$
@@
    Negation Symmetry: The weights remain unchanged under the transformation $z\mapsto-z$. In other words, they are not biased toward positive or negative solution values. To ensure this property, a sufficient condition is to choose $C_{1},C_{2}$ to be functions of features invariant to this negation transformation. For example,
@@
3.2 SP-WENOc [44]
While SP-WENO leads to an entropy stable scheme, numerical results presented in [44] (also see Section 5) demonstrate its poor performance near discontinuities in the TeCNO framework. As discussed in [44], this issue can be attributed to the fact that the reconstructed jump $\llbracket z\rrbracket_{i+\frac{1}{2}}$ is zero in most cases, resulting in zero numerical diffusion (see (13)) in these regions. While the absence of numerical diffusion may be acceptable in smooth regions, this can cause Gibbs oscillations near discontinuities. The most important problematic cases lie in the so-called C-region [44], where the jump ratios are either $\theta_{i}^{+}<1$, $\theta_{i+1}^{-}>1$ or $\theta_{i}^{+}>1$, $\theta_{i+1}^{-}<1$ (cases 2 and 3 as discussed in Section 3.3).
@@ 3.2 SP-WENOc [44]
To keep the weights consistent, the following additional modification is introduced
$C_{1}^{\#}=\min{\left(\max{\left(\overline{C}_{1},-\frac{3}{8}\right)},\frac{1}{8}\right)},\hskip 8.53581ptC_{2}^{\#}=\min{\left(\max{\left(\overline{C}_{2},-\frac{3}{8}\right)},\frac{1}{8}\right)}.$ (25)
@@ 3.3 Feasible Region for SP-WENO
The SP-WENO and SP-WENOc formulations are just two possible weight selection strategies satisfying the constraints that guarantee consistency, sign-preservation, and third-order accuracy. Note that none of these constraints describe the behavior of solutions near shocks, and so these two reconstruction methods are not necessarily constructed with precise shock-capturing in mind. That is to say, there are potentially other SP-WENO variants that perform better near discontinuities while satisfying the constraints.
As noted in [16], the reconstruction problem at an interface can be broken down in several cases depending on the jump ratios $\theta_{i}^{+},\theta_{i+1}^{-}$. Each case has its own constraints on the perturbations $(C_{1},C_{2})$ such that consistency, sign-preservation, and third-order accuracy are guaranteed. We first define the notion of a feasible region based on the values of $\theta_{i}^{+},\theta_{i+1}^{-}$.
@@
    Sign Property: The choice of $(C_{1},C_{2})$ lead to weights that satisfy the constraint (18).
@@
In Table 1, we list the various cases of interest along with the corresponding sign property and accuracy constraints. Note that the various $\Omega_{\Theta}$ listed in the table (column 2) form a disjoint partition of ${\mathbb{R}}^{2}$. Further, $\psi^{+}_{i+\frac{1}{2}}$ is used to describe $\Omega_{\Theta}$ in cases 2 and 3, which is a function of $\theta^{+}_{i},\theta^{-}_{i+1}$ by virtue of (16). We reintroduce the following notation from [16] (with the $i+\frac{1}{2}$ subscript dropped from $\psi^{+},\psi^{-}$ terms)
:=$\mathcal{L}:=\begin{cases}\frac{C_{1}}{\frac{1}{8}(1+\psi^{+})}+\frac{C_{2}}{\frac{1}{8}(1+\psi^{-})},&\text{ if }\psi^{+}\neq-1\\ C_{1}-C_{2}+1,&\text{ if }\psi^{+}=\psi^{-}=-1\end{cases},$ (26)
which is also used to define the sign property constraints in Table 1. The expression of $\mathcal{L}$ can be obtained from the constraint (18) through simple algebraic manipulations. For full details, we refer interested readers to [16].
The SP-WENO and SP-WENOc formulations satisfy the constraints of each scenario. As will be discussed in the DL-based DSP-WENO formulation, we carefully consider these scenarios while constructing the reconstruction method learned via a neural network.

### v2

@@ 3 Sign-Preserving WENO Reconstructions
A typical WENO reconstruction [jiangshu96] produces a $(2k-1)$th-order accurate reconstruction by taking a convex combination of the $12k-$ 1 candidate polynomials considered in the $k$th-order ENO reconstruction. A third-order WENO has the following form for the left and right reconstructions of the variable $z$ at the interface $x_{i+\frac{1}{2}}$:
$z_{i+\frac{1}{2}}^{-}{=}\frac{w_{0}(z_{i}{+}z_{i+1})+w_{1}(3z_{i}{-}z_{i-1})}{2},\ z_{i+\frac{1}{2}}^{+}{=}\frac{\tilde{w}_{0}(3z_{i+1}{-}z_{i+2})+\tilde{w}_{1}(z_{i}{+}z_{i+1})}{2},$ (12)
where the weights $w_{0},w_{1},\tilde{w}_{0},\tilde{w}_{1}$ (at each interface) must be chosen to: (i) ensure third-order accuracy of the reconstructions $z_{i+\frac{1}{2}}^{\pm}$ for smooth solutions and (ii) provide minimal weight to linear polynomials on stencils containing discontinuities. In general, WENO reconstructions do not satisfy the sign property, and their use in the TeCNO framework described by (11) does not yield entropy stable schemes. Thus, a WENO reconstruction must be specifically designed to be sign-preserving, i.e., it must satisfy the sign property (10).
3.1 SP-WENO [fjordholm2016sign]
@@
$C_{1}(\theta_{i}^{+},\theta_{i+1}^{-})=\begin{cases}\frac{1}{8}\left(\frac{\kappa^{+}}{(\kappa^{+})^{2}+(\kappa^{-})^{2}}\right)&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}<0,\psi^{+}\neq-1\\ 0&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}=-1\\ -\frac{3}{8}&\text{ if }\theta_{i}^{+}=1\text{ or }\psi^{+}\geq 0,|\theta_{i}^{+}|\leq 1\\ \frac{1}{8}&\text{ if }\psi^{+}\geq 0,|\theta_{i}^{+}|>1\end{cases},$
@@
$\kappa^{+}(\theta_{i}^{+},\theta_{i+1}^{-}):=\begin{cases}\frac{1}{1+\psi^{+}}&\text{ if }\theta_{i}^{+}\neq 1,\psi^{+}\neq-1\\ 1&\text{ otherwise}\end{cases},\hskip 14.22636pt\kappa^{-}(\theta_{i}^{+},\theta_{i+1}^{-}):=\kappa^{+}(\theta_{i+1}^{-},\theta_{i}^{+}).$
@@
    Negation Symmetry: The weights remain unchanged under the transformation $z\mapsto-z$. A sufficient condition to ensure this property is to choose $C_{1},C_{2}$ to be functions of features invariant to this negation transformation. For example,
@@
3.2 SP-WENOc [ray2018third]
While SP-WENO leads to an entropy stable scheme, numerical results presented in [ray2018third] (also see Section 5) demonstrate its poor performance near discontinuities in the TeCNO framework. As discussed in [ray2018third], this issue can be attributed to the fact that the reconstructed jump $\llbracket z\rrbracket_{i+\frac{1}{2}}$ is zero in most cases, resulting in zero numerical diffusion (see (11)) in these regions. While the absence of numerical diffusion may be acceptable in smooth regions, this can cause Gibbs oscillations near discontinuities. The most important problematic cases lie in the so-called C-region [ray2018third], where the jump ratios are either $\theta_{i}^{+}<1$, $\theta_{i+1}^{-}>1$ or $\theta_{i}^{+}>1$, $\theta_{i+1}^{-}<1$.
@@ 3.3 Feasible Region for SP-WENO
The SP-WENO and SP-WENOc formulations are just two possible weight selection strategies satisfying the constraints that guarantee consistency, sign-preservation, and third-order accuracy. Note that none of these constraints describe the behavior of solutions near shocks, thus SP-WENO and SP-WENOc are not necessarily constructed with precise shock-capturing in mind. That is to say, there are potentially other SP-WENO variants that perform better near discontinuities while satisfying the constraints.
As noted in [fjordholm2016sign], the reconstruction problem at an interface can be broken down in several cases depending on the jump ratios $\theta_{i}^{+},\theta_{i+1}^{-}$. Each case has its own constraints on the perturbations $(C_{1},C_{2})$ such that consistency, sign-preservation, and third-order accuracy are guaranteed. We define the notion of a feasible region based on the values of $\theta_{i}^{+},\theta_{i+1}^{-}$.
@@
    Sign Property: The choice of $(C_{1},C_{2})$ leads to weights that satisfy the constraint (16).
@@
Thus, for any given values of $\theta_{i}^{+},\theta_{i+1}^{-}$, a feasible region is defined. See Section SM1 in the Supplementary Material (SM) for further discussion on the cases and their associated feasible regions.
4 DSP-WENO
To this end, an artificial neural network ${\bm{\mathcal{N}}}_{{\bm{\Phi}}}:\mathbb{R}^{N_{I}}\mapsto\mathbb{R}^{N_{O}}$ with trainable parameters $\bm{\Phi}$ (known as weights and biases) is trained on a suitable dataset such that an objective/loss function $\Pi(\bm{\Phi})$ is optimized.
For the $k$-th sample in $\mathbb{T}$, $\bm{X}^{(k)}\in{\mathbb{R}}^{4}$ consists of the four cell center values of some function $z$ in the local stencil centered about (some) $x_{i+\frac{1}{2}}$: $[z_{i-1},z_{i},z_{i+1},z_{i+2}]$, while $\bm{Y}^{(k)}\in{\mathbb{R}}^{2}$ is the true cell interface values $z^{\pm}(x_{i+\frac{1}{2}})$. Note that the interface values from the left (-) and right (+) will be identical for a smooth function, but will be different if there is a discontinuity at the interface.
We want to ensure that all the inputs to the neural network are of the same order of magnitude (essential for generalization). Thus, we first scale the cell center values as
A similar input scaling was also considered in [ray2018artificial, ray2019detecting]. Using the scaled cell center values $[z_{i-1}^{*},z_{i}^{*},z_{i+1}^{*},z_{i+2}^{*}]$, we compute $\theta_{i+1}^{-},\theta_{i}^{+},\left|\Delta z_{i-\frac{1}{2}}^{*}\right|,\left|\Delta z_{i+\frac{1}{2}}^{*}\right|,\left|\Delta z_{i+\frac{3}{2}}^{*}\right|$. We note that the jump ratios $\theta_{i+1}^{-},\theta_{i}^{+}$ are invariant to the scaling (24). Thus we define the map ${\bm{\mathcal{S}}}:{\mathbb{R}}^{4}\rightarrow{\mathbb{R}}^{5}$ which transforms $\bm{X}^{(k)}$ in the training set to the jump ratios and scaled absolute jumps mentioned above, i.e.,
[…]
The output of the network ${\bm{\mathcal{N}}}_{\bm{\Phi}}$ is the vector $\bm{\alpha}\in{\mathbb{R}}^{5}$ whose components satisfy
[…]
Let us define a function ${\bm{\mathcal{V}}}:{\mathbb{R}}^{5}\rightarrow{\mathbb{R}}^{2\times 5}$ that maps ${\bm{\mathcal{S}}}(\bm{X}^{(k)})$ to 5 vertices in ${\mathbb{R}}^{2}$, i.e.,
In (28), the $\bm{\nu}^{(k)}_{s}$ corresponds to the vertices of a convex polygon (defining the feasible region), thus ensuring that any convex combination of these vertices will result in a vector in ${\mathbb{R}}^{2}$ that lies in the (closure) of this convex polygon. These convex regions are formed by at most five vertices. In situations with less than five vertices, some of the $\bm{\nu}^{(k)}_{s}$ in (28) are replaced by a repeated interior node of the convex region, typically the centroid. Further details about ${\bm{\mathcal{V}}}$ can be found in Section 4.1. The output of ${\bm{\mathcal{N}}}_{\bm{\Phi}}$ is combined with these vertices to obtain the DSP-WENO weight perturbations $C_{1},C_{2}$, given by
[…]
We finally define the objective function in terms of the trainable parameters ${\bm{\Phi}}$ of the network using a mean squared error (MSE) loss
$\Pi(\bm{\Phi})$ $=\frac{1}{K}\sum_{k=1}^{K}\Pi_{k}({\bm{\Phi}}),\quad\Pi_{k}({\bm{\Phi}})=\left\|\widehat{\bm{Y}}^{(k)}-\bm{Y}^{(k)}\right\|_{2}$ (31)
where the $\widehat{\bm{Y}}^{(k)}$ encapsulate the dependence on $\bm{\Phi}$ and .$\|.\|_{2}$ denotes the Euclidean 2-norm. Thus, the network is trained by solving the minimization problem
${\bm{\Phi}}^{\dagger}=\underset{{\bm{\Phi}}}{\arg\min}\Pi(\bm{\Phi}).$